QUESTION 11 An advance in technology in the production of good X causes a. a leftward shift in the supply curve for good X. bQUESTION 13 -- you incur an Every time you make a a. competitive move; price b. decision; efficient effect c. choice; opportuQUESTION 18 If Cassandra bought 12 blouses last year when her income was $46,000 and she buys 14 blouses this year when her iif demand is If the seller of good X raises the price of good X. it follows that the total revenue of good X will a rise; elaGood Z is income unit elastic. This means that the percentage change in income is a. equal to the percentage change in the qu
Answer #1

11. Ans: a rightward shift in the supply curve of good X

Explanation:

Due to advancement in technology, the production and supply of good X will increase. So, the supply curve of good X will shift rightward.

13. Ans: choice ; opportunity cost

18. Ans: +1.26

Explanation:

Here, Y1 = 46,000     Q1 = 12

          Y2 = 52,000     Q2 = 14

Income elasticity of demand = (Q2 - Q1) / (Y2 - Y1) * (Y1 + Y2) / (Q1 + Q2)

                                             = (14 - 12) / (52,000 - 46,000) * (46,000 + 52,000) / (12 + 14)

                                             = (2 / 6,000) * (98,000 / 26)

                                             = 196,000 / 156,000

                                             = +1.26

Ans: fall ; elastic

Explanation:

If demand is elastic, a 1% increase in price lead to more than 1% decrease in quantity demanded. So, the price rise in price leads to decrease in quantity demanded.

Ans: equal to the percentage change in the quantity demanded of good Z.